# InstructPipe: Building Visual Programming Pipelines with Human Instructions Using LLMs

### v2

@@ 7.3. Instructing LLMs Poses Challenges for Both Novices and, Potentially, Experts
As we discussed above, non-experts found it challenging to instruct LLM. More interestingly, we found that even we, the inventors of InstructPipe, failed to write optimal instructions. For instance, the two captions of 13(c) are “Describe the image and turn it into a cat image” and “Edit an image by updating the image caption”. Neither caption explicitly describes the detailed pipeline flow clearly, and therefore, all the six evaluation trials (section 5) were incomplete (see 9(a) for one example). The average ratio of user interactions is $45.8\%$, more than twice the average value for our multimodal pipelines ($20.8\%$). To further understand the cause of the failure, another author improved the instruction into “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”.
@@
The resulting pipeline is significantly improved but still not perfect (9(b)). The user only needs to turn “Imagen” into another mode so that it also accepts the input “image” node. Revisiting the improved instruction, we instructed InstructPipe with “generate a cat image based on THE updated caption”, which actually missed the input image.
The important takeaway is while natural languages are proven to be one promising communication media that connects humans and AI systems (Chen et al., 2023; Wang et al., 2023b), instructions may not be the best format to facilitate such connection. We believe the reason is that instructions are still not intuitive to humans: AI typically requires flawless and unambiguous instructions while humans tend to express their intention using ambiguous natural languages in conversations. We encourage future work to investigate alternative interaction mediums beyond instructions to further enhance user experience in human-AI collaboration.
@@ 8.1. Assisting Humans to Prompt AI Copilot in Visual Programming
InstructPipe introduces a novel user interaction technique for visual programming, along with its set of challenges – prompting AI is not easy (Zamfirescu-Pereira et al., 2023). While latest research has explored prompt writing assistants (Brade et al., 2023; LangChain, 2024), creating such assistants in visual programming poses unique challenges, as discussed in section 7.2, and requires further dedicated investigation. Despite these challenges, the visual programming workspace offers unique opportunity – it provides an interactive and visual medium for delivering AI-generated information. For example, a prompt writing assistant could provide “a pipeline preview” in real time via a lightweight LLM. Visualizing estimated outcomes, such as unexpected pipeline results (as illustrated in 9(a)), highlights the potential of these tools to guide users in refining their instructions effectively.
@@ 8.2. Node Parameter Tuning
InstructPipe focuses on generating the graph structure in the pipeline (section 4.1), and InstructPipe is not able to generate node parameters. Latest research in AI agents show great potentials of distributing a systematic task among multiple LLMs and let them solve the problem collaboratively (Jin et al., 2024). We encourage future work to extend such distributed AI agent approaches to generate suitable node parameters to further reduce users’ workload in tuning them.
@@ 8.3. A Larger and Dynamic Node Library
InstructPipe is a AI assistant prototype on a small scale library with 27 nodes. Similar to other tool-calling LLM systems (Schick et al., 2024; De La Torre et al., 2024), InstructPipe cannot generate any out-of-scope node, and thus there is a limited scope of pipelines that InstructPipe can generate. Future work should investigate a scale-up problem by creating an assistant that support a large-scale nodes (Qin et al., 2023a). What new technical challenge will emerge? Will such a large-scale library provide practical human value? If yes, what are the concrete new user experience it opens up in visual programming?
Additionally, future work should explore a dynamic solution of the node library, in which an AI assistant can define necessary nodes in visual programming on the fly. HuggingGPT (Shen et al., 2024) is a pioneering project that shares similar vision as this goal, but existing investigations show that the accuracy of such open-ended generation is still unsatisfactory (Patil et al., 2023; Qin et al., 2023b). How can we design an interface to bridge such imperfect AI and human users in visual programming copilot? What will be the interaction paradigm in an interactive system with a dynamic node library?
@@ 8.4. Refining System Component Design
Prompt Design. We leveraged in-context learning capability of LLMs in our prompt design. Prior work shows that few-shot examples cause bias effects in practice (Zhao et al., 2021b), and thus we encourage future work to mitigate this bias by collecting a large dataset and finetuning LLM on the dataset.
@@ 8.5. Evaluation Metrics and Long-term Evaluation
In the user evaluation, we conducted a lab study to understand the user experience of InstructPipe. As future work, we plan to conduct longer-term studies and gather more realistic insights than those we obtained from the lab study. This is critical for us to understand the long-term usefulness of our assistant for beginners, as well as collecting feedback to inform our system design.
@@ B.1.2. Code Writer
The prompt in Code Writer is dynamic, which is dependent on the nodes chosen in Node Selector. Therefore, we cannot provide all the possible prompts in the supplementary materials. Here, we will focus on providing examples of two detailed node configurations utilized in InstructPipe. Fig. 6 shows the structure of the prompt utilized in this LLM stage. Fig. 10 provides two examples of node configurations (i.e., “Body segmentation” and “PaLI”) that InstructPipe may chose into the highlighted line(s). Each configuration includes keys of “nodeSpecId” (i.e., node types), “description”, “category” and “examples”. For those nodes that support input and output edges, “inputSpecs” and “outputSpecs” specify the sockets and their corresponding valid data types. For example, the output socket name of “Body segmentation” is “segmentationResult”, and its data type is “masks”. Some nodes (e.g., “Body segmentation”) include recommended node(s) (e.g., “Mask visualizer” for “Body segmentation”), and our configuration also contains such information in the dictionary.
@@ B.2. Code Interpreter
Here we provide more low-level implementation details on Code Interpreter. The Code Interpreter parses generated pseudocode into a visual programming pipeline for visualization at the Visual Blocks workspace. Fig. 12 shows the data type definition of graphs, nodes, and edges in the system. The example JSON file to be parsed into the Typescript interface is available at the Visual Blocks website [^5] https://visualblocks.withgoogle.com/. The JSON file is available for data structure exploration by 1) entering an example project and 2) clicking on the “Export” button on the top-right corner.. Our code defines a visual programming pipeline into an array of serialized nodes, $G(N)$. When the user adds a new node to an pipeline, it is adding a new “SerializedNode”, where it contains the edge definition between this new node and other nodes in the current workspace, to the current “SerialedGraph”. This mechanism ensures that nodes can be incrementally added in the order they appear in the pseudocode order while maintaining the integrity of the graph by clearly defining dependencies and data flow between nodes. Algorithm 1 further shows how InstructPipe parses code and incrementally add nodes to formulate a final serialized graph.
@@
[ framesep=2mm, baselinestretch=1.2, bgcolor=LightGray, fontsize=, ] typescript /** A serialized graph. */ export declare interface SerializedGraph nodes: SerializedNode[];
@@
/** Serialized incoming edges * that connect to this node. */ incomingEdges?: [inputId: string]: SerializedIncomingEdge[] ;
@@
Figure 12. The definition of a graph, a node and an edge in the system using the Typescript language. Only core properties of graphic structure definition are presented in this figure.
@@
3 Output: $G(N)$: a visual programming pipeline ($SerializedGraph$) that mainly stores an array of $SeirializedNode$ (Fig. 12).
@@ 5       $T=tokenizer(line)$
$G=optim\_layout(G)$ // Perform the UI layout optimization, as shown in Fig. 11
@@
Appendix C Technical Evaluation
((a)) Search news from Google, summarize it and then conduct fact check. Input: a keyword for Google Search; Output: a summarization of the news and a fact-check result.
@@ C.1. Data Post-Processing
The observation motivated us to post-process the raw data to present more rigorous evaluation results. We first removed incomplete pipelines and the isolated graphs in each pipeline (if there are any). To further enhance the annotation quality, two authors individually annotated the caption of each pipeline separately by referring to the original captions and pipelines authored by the participants. It is important to note that we finished the workshop and the data annotation task before we completed the system implementation. The two authors had no experience using InstructPipe before completing the annotation. We believed this process could effectively enhance the quality of the captions while maintain fairness of the technical evaluation.
@@
We removed the pipelines that contain node(s) out of our focus 27 nodes, and selected all the remaining pipelines as our final evaluation set. We argue that this post-processing is critical for reporting a fair accuracy value since InstructPipe can only generate pipelines based on its known node library. The final 48 pipelines (out of 64 pipelines) are comprised of 23 language pipelines, seven visual pipelines, and 18 multi-modal pipelines. Fig. 13 shows three pipelines created by the participants. 13(b) is an example of the pipelines that include out-of-scope nodes, and therefore is not included in the final 48 pipelines. In the technical evaluation, we ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions from two authors for each pipelines) and evaluated the generation results using the metric that will be introduced below.
@@ D.2. User Study Pipelines
Figure 15. Real-time multimodal pipeline. The “Keyword to image” node is used to search sunglasses image, and the “Virtual sticker” node anchors the sunglasses onto users’ face.
Fig. 14 and Fig. 15 visualize two pipelines we required the participants to complete in our user study. Fig. 15 is a multimodal pipeline that allows participants to interact with AR effects in real time. Our technical evaluation shows that InstructPipe can generate this pipeline accurately: the averaged ratio of human interactions = 5.2%. Fig. 15 is a text-based pipeline that provides participants with a summary of the news searched from Google. The technical evaluation reveals that InstructPipe cannot generate this pipeline accurately without further human interaction, and the averaged ratio of additional human interactions is 27.8%. While the generated diagram (with error) is not deterministic, we observed that InstructPipe commonly generates the pipeline in Figure 14 without “URL to HTML” or “PaLM Text Generator” nodes. The error implies that the LLM may misinterpret 1) the data from the “selected text” port of the “String picker” node is the texts on the web instead of the web URL and 2) that “Text processor” has the LLM capability to process the texts instead of simply combining two texts.
Note that even though InstructPipe may be able to complete the pipeline structure in Fig. 15 from users’ instruction, we observed that participants still need to fine-tune their keywords to get an ideal pair of sunglasses. Additionally, the default anchor value is “Face top”, so participants need to use the drop-down menu on the “Virtual sticker” node to change the value to “Eyes”. This further motivates us to use the metric of “Time” in addition to the number of user interactions in our study. Our demo video also covers the workflows of these two pipelines.
@@ D.3. Assistant Provided to the Participants in the User Evaluation
In the user evaluation, our goal is to make the interface condition (either InstructPipe or Visual Blocks) as the only independent variable that changes our dependent variables (section 5.2). Similar to user evaluations of other early-stage HCI research, we had to improvise for lacking system affordances. As an example, we would include help menus and error recovery modals in the future versions of our system but at this early stage we relied on in-person help to nudge and assist our user study participants. We took actions (i.e., assistants) in the user evaluation to ensure the study is under an appropriate amount of control as well as maintain the fairness of our study.
Here we elaborate two examples of assistants we provided in the user study.
@@
In the InstructPipe condition, one participant started their “instructions” by dragging a text box into the visual programming workspace and began typing. When noticing this issue, we kindly asked the participant whether s/he wanted to write instructions or build a pipeline from scratch. S/he then noticed this issue and clicked on the “InstructPipe” button to write prompts. Note that we explicitly taught every participant how to use InstructPipe and asked participants themselves to go through the instruction processes in the training task (Fig. 7).
@@
In the Visual Blocks condition, one participant first dragged a “Virtual sticker” into the workspace when s/he wants to build the multimodal pipeline as required (Fig. 15). After a while, s/he asked us for the meaning of “landmarks” on the first input port of the “Virtual sticker” node (Fig. 15). We then answered this question and provided a hint on the “Face landmark” node (Fig. 15) that could produce the “Face landmarks” required by the “Virtual sticker”. While we had explained all the nodes that the participants need to use in the study in the training task (Fig. 7), we consider such technical questions as reasonable because all of our participants are non-experts. Programming itself is a difficult skill, and it is quite common that people may forget some of the knowledge that they have just learned. Instead of being silent and keeping the participants stuck on a technical issue, we believed offering technical help is an important action we must take to ensure the data quality we collected in the study.
@@
These anecdotes in the user evaluation reveal several limitations of the visual programming system: some designs may not be very intuitive to non-experts. Since the goal of our user evaluation is understanding the benefits of InstructPipe compared to Visual Blocks (without AI assistants), we made our best efforts to take actions to prevent the effects caused by other factors from influencing our data. Meanwhile, we also encourage future work to further explore the system design so that future users can more easily use our assistant in visual programming.
@@
Table 4. The counterbalance sheet of the user evaluation. Each cell is in the format of “Interface / Pipeline”. “Instruct” and “VB” mean the “InstructPipe” and “Visual Blocks” conditions, respectively. “Search” and “Tryon” represent the “text-based pipeline” (Fig. 14) and the “real-time multimodal pipeline” (Fig. 15), respectively.

### v3

@@ 7.3. Instructing LLMs Poses Challenges for Both Novices and, Potentially, Experts
As we discussed above, non-experts found it challenging to instruct LLM. More interestingly, we found that even we, the inventors of InstructPipe, failed to write optimal instructions. For instance, the two captions of 13(c) are “Describe the image and turn it into a cat image” and “Edit an image by updating the image caption”. Neither caption explicitly describes the detailed pipeline flow clearly, and therefore, all the six evaluation trials (section 5) were incomplete (see 9(a) for one example). The average ratio of user interactions is $45.8\%$, more than twice the average value for our multimodal pipelines ($20.8\%$). To further understand the cause of the failure, another author improved the instruction into “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”. The resulting pipeline is significantly improved but still not perfect (9(b)). The user only needs to turn “Imagen” into another mode so that it also accepts the input “image” node. Revisiting the improved instruction, we instructed InstructPipe with “generate a cat image based on THE updated caption”, which actually missed the input image.
The important takeaway is while natural languages are proven to be one promising communication media that connects humans and AI systems (Chen et al., 2023; Wang et al., 2023b), instructions may not be the best format to facilitate such connection. We believe the reason is that instructions are still not intuitive to humans: AI typically requires flawless and unambiguous instructions, while humans tend to express their intentions using ambiguous natural languages in conversations. We encourage future work to investigate alternative interaction mediums beyond instructions to further enhance user experience in human-AI collaboration.
@@ 8.1. Assisting Humans to Prompt AI Copilot in Visual Programming
InstructPipe introduces a novel user interaction technique for visual programming, along with its set of challenges – prompting AI is not easy (Zamfirescu-Pereira et al., 2023). While the latest research has explored prompt writing assistants (Brade et al., 2023; LangChain, 2024; Ma et al., 2024), creating such assistants in visual programming poses unique challenges, as discussed in section 7.2, and requires further dedicated investigation. Despite these challenges, the visual programming workspace offers a unique opportunity – it provides an interactive and visual medium for delivering AI-generated information. For example, a prompt writing assistant could provide “a pipeline preview” in real time via a lightweight LLM. Visualizing estimated outcomes, such as unexpected pipeline results (as illustrated in 9(a)), highlights the potential of these tools to guide users in refining their instructions effectively.
@@ 8.2. Node Parameter Tuning
InstructPipe focuses on generating the graph structure in the pipeline (section 4.1), and InstructPipe is not able to generate node parameters. The latest research in AI agents shows great potential for distributing a systematic task among multiple LLMs and letting them solve the problem collaboratively (Jin et al., 2024). We encourage future work to extend such distributed AI agent approaches to generate suitable node parameters to further reduce users’ workload in tuning them.
@@ 8.3. A Larger and Dynamic Node Library
InstructPipe is an AI assistant prototype on a small-scale library with 27 nodes. Similar to other tool-calling LLM systems (Schick et al., 2024; De La Torre et al., 2024), InstructPipe cannot generate any out-of-scope node, and thus, there is a limited scope of pipelines that InstructPipe can generate. Future work should investigate a scale-up problem by creating an assistant that supports large-scale nodes (Qin et al., 2023a). What new technical challenge will emerge? Will such a large-scale library provide practical human value? If yes, what are the concrete new user experience it opens up in visual programming?
Additionally, future work should explore a dynamic solution of the node library, in which an AI assistant can define necessary nodes in visual programming on the fly. HuggingGPT (Shen et al., 2024) is a pioneering project that shares a similar vision as this goal, but existing investigations show that the accuracy of such open-ended generation is still unsatisfactory (Patil et al., 2023; Qin et al., 2023b). How can we design an interface to bridge such imperfect AI and human users in visual programming copilot? What will be the interaction paradigm in an interactive system with a dynamic node library?
@@ 8.4. Refining System Component Design
Prompt Design. We leveraged the in-context learning capability of LLMs in our prompt design. Prior work shows that few-shot examples cause bias effects in practice (Zhao et al., 2021b), and thus, we encourage future work to mitigate this bias by collecting a large dataset and finetuning LLM on the dataset.
@@ 8.5. Evaluation Metrics and Long-term Evaluation
In the user evaluation, we conducted a lab study to understand the user experience of InstructPipe. As future work, we plan to conduct longer-term studies and gather more realistic insights than those we obtained from the lab study. This is critical for us to understand the long-term usefulness of our assistant for beginners, as well as collect feedback to inform our system design.
@@ 9. Conclusion
Acknowledgements.
We would like to thank Gong Xuan, Fengyuan Zhu, Kevin Zhang and Karl Rosenberg for their feedback and discussion on our early-stage prototype, as well as Koji Yatani and Takeo Igarashi for the feedback on the paper draft. We also thank our reviewers for their insightful feedback.
@@ B.1.2. Code Writer
The prompt in Code Writer is dynamic, which is dependent on the nodes chosen in Node Selector. Therefore, we cannot provide all the possible prompts in the supplementary materials. Here, we will focus on providing examples of two detailed node configurations utilized in InstructPipe. Figure 6 shows the structure of the prompt utilized in this LLM stage. Figure 10 provides two examples of node configurations (i.e., “Body segmentation” and “PaLI”) that InstructPipe may chose into the highlighted line(s). Each configuration includes keys of “nodeSpecId” (i.e., node types), “description”, “category” and “examples”. For those nodes that support input and output edges, “inputSpecs” and “outputSpecs” specify the sockets and their corresponding valid data types. For example, the output socket name of “Body segmentation” is “segmentationResult”, and its data type is “masks”. Some nodes (e.g., “Body segmentation”) include recommended node(s) (e.g., “Mask visualizer” for “Body segmentation”), and our configuration also contains such information in the dictionary.
@@ B.2. Code Interpreter
Here, we provide more low-level implementation details on Code Interpreter. The Code Interpreter parses generated pseudocode into a visual programming pipeline for visualization at the Visual Blocks workspace. Figure 12 shows the data type definition of graphs, nodes, and edges in the system. The example JSON file to be parsed into the Typescript interface is available at the Visual Blocks website [^5] https://visualblocks.withgoogle.com/. The JSON file is available for data structure exploration by 1) entering an example project and 2) clicking on the “Export” button on the top-right corner.. Our code defines a visual programming pipeline into an array of serialized nodes, $G(N)$. When the user adds a new node to a pipeline, it adds a new “SerializedNode”, containing the edge definition between this new node and other nodes in the current workspace, to the current “SerialedGraph”. This mechanism ensures that nodes can be incrementally added in the order they appear in the pseudocode order while maintaining the integrity of the graph by clearly defining dependencies and data flow between nodes. Algorithm 1 further shows how InstructPipe parses code and incrementally adds nodes to formulate a final serialized graph.
@@
[ framesep=2mm, baselinestretch=1, bgcolor=LightGray, fontsize=, ] typescript /** A serialized graph. */ export declare interface SerializedGraph nodes: SerializedNode[];
@@
/** * Serialized incoming edges that * connect to this node. */ incomingEdges?: [inputId: string]: SerializedIncomingEdge[] ;
@@
Figure 12. The definition of a graph, a node and an edge in the system using the Typescript language. Only the core properties of graphic structure definition are presented in this figure.
@@
3 Output: $G(N)$: a visual programming pipeline ($SerializedGraph$) that mainly stores an array of $SeirializedNode$ (Figure 12).
@@ 5       $T=tokenizer(line)$
$G=optim\_layout(G)$ // Perform the UI layout optimization, as shown in Figure 11
@@
((a)) Search news from Google, summarize it, and then conduct a fact check. Input: a keyword for Google Search; Output: a summarization of the news and a fact-check result.
@@ C.1. Data Post-Processing
The observation motivated us to post-process the raw data to present more rigorous evaluation results. We first removed incomplete pipelines and the isolated graphs in each pipeline (if there are any).
To further enhance the annotation quality, two authors individually annotated the caption of each pipeline separately by referring to the original captions and pipelines authored by the participants. It is important to note that we finished the workshop and the data annotation task before we completed the system implementation. The two authors had no experience using InstructPipe before completing the annotation. We believed this process could effectively enhance the quality of the captions while maintaining the fairness of the technical evaluation.
@@
We removed the pipelines that contain node(s) out of our focus 27 nodes, and selected all the remaining pipelines as our final evaluation set. We argue that this post-processing is critical for reporting a fair accuracy value since InstructPipe can only generate pipelines based on its known node library. The final 48 pipelines (out of 64 pipelines) are comprised of 23 language pipelines, seven visual pipelines, and 18 multi-modal pipelines. Figure 13 shows three pipelines created by the participants. 13(b) is an example of the pipelines that include out-of-scope nodes, and therefore are not included in the final 48 pipelines. In the technical evaluation, we ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions from two authors for each pipeline) and evaluated the generation results using the metric that will be introduced below.
@@ D.2. User Study Pipelines
Figure 15. Real-time multimodal pipeline. The “Keyword to image” node is used to search a sunglasses image, and the “Virtual sticker” node anchors the sunglasses onto the user’ face.
Figure 14 and Figure 15 visualize two pipelines we required the participants to complete in our user study. Figure 15 is a multimodal pipeline that allows participants to interact with AR effects in real time. Our technical evaluation shows that InstructPipe can generate this pipeline accurately: the averaged ratio of human interactions = 5.2%. Figure 15 is a text-based pipeline that provides participants with a summary of the news searched from Google. The technical evaluation reveals that InstructPipe cannot generate this pipeline accurately without further human interaction, and the average ratio of additional human interactions is 27.8%. While the generated diagram (with error) is not deterministic, we observed that InstructPipe commonly generates the pipeline in Figure 14 without “URL to HTML” or “PaLM Text Generator” nodes. The error implies that the LLM may misinterpret 1) the data from the “selected text” port of the “String picker” node is the texts on the web instead of the web URL and 2) that “Text processor” has the LLM capability to process the texts instead of simply combining two texts.
Note that even though InstructPipe may be able to complete the pipeline structure in Figure 15 from users’ instruction, we observed that participants still need to fine-tune their keywords to get an ideal pair of sunglasses. Additionally, the default anchor value is “Face top”, so participants need to use the drop-down menu on the “Virtual sticker” node to change the value to “Eyes”. This further motivates us to use the metric of “Time” in addition to the number of user interactions in our study. Our demo video also covers the workflows of these two pipelines.
@@ D.3. Assistant Provided to the Participants in the User Evaluation
In the user evaluation, our goal is to make the interface condition (either InstructPipe or Visual Blocks) as the only independent variable that changes our dependent variables (section 5.2). Similar to user evaluations of other early-stage HCI research, we had to improvise for lacking system affordances. As an example, we would include help menus and error recovery models in the future versions of our system, but at this early stage, we relied on in-person help to nudge and assist our user study participants. We took actions (i.e., assistants) in the user evaluation to ensure the study is under an appropriate amount of control as well as maintain the fairness of our study.
Here, we elaborate on two examples of assistants we provided in the user study.
@@
In the InstructPipe condition, one participant started their “instructions” by dragging a text box into the visual programming workspace and began typing. When noticing this issue, we kindly asked the participant whether s/he wanted to write instructions or build a pipeline from scratch. S/he then noticed this issue and clicked on the “InstructPipe” button to write prompts. Note that we explicitly taught every participant how to use InstructPipe and asked participants themselves to go through the instruction processes in the training task (Figure 7).
@@
In the Visual Blocks condition, one participant first dragged a “Virtual sticker” into the workspace when s/he wanted to build the multimodal pipeline as required (Figure 15). After a while, s/he asked us for the meaning of “landmarks” on the first input port of the “Virtual sticker” node (Figure 15). We then answered this question and provided a hint on the “Face landmark” node (Figure 15) that could produce the “Face landmarks” required by the “Virtual sticker”. While we had explained all the nodes that the participants need to use in the study in the training task (Figure 7), we consider such technical questions reasonable because all of our participants are non-experts. Programming itself is a difficult skill, and it is quite common that people may forget some of the knowledge that they have just learned. Instead of being silent and keeping the participants stuck on a technical issue, we believed offering technical help was an important action we must take to ensure the data quality we collected in the study.
@@
These anecdotes in the user evaluation reveal several limitations of the visual programming system: some designs may not be very intuitive to non-experts. Since the goal of our user evaluation is understanding the benefits of InstructPipe compared to Visual Blocks (without AI assistants), we made our best efforts to take action to prevent the effects caused by other factors from influencing our data. Meanwhile, we also encourage future work to further explore the system design so that future users can more easily use our assistant in visual programming.
@@
Table 4. The counterbalance sheet of the user evaluation. Each cell is in the format of “Interface / Pipeline”. “Instruct” and “VB” mean the “InstructPipe” and “Visual Blocks” conditions, respectively. “Search” and “Tryon” represent the “text-based pipeline” (Figure 14) and the “real-time multimodal pipeline” (Figure 15), respectively.